OpenStudy (honeymonster):

please help me solve this simultaneous equation.. x+2y=5 3x+2y=3

OpenStudy (anonymous):

Subtract them, what will you get?

OpenStudy (honeymonster):

subtract what?

OpenStudy (anonymous):

x+2y=5 (1) 3x+2y=3 (2) Subtract the first equation for the second one, of course!

OpenStudy (honeymonster):

i'm sorry i've never been taught this in school..its difficult for me to understand :/

OpenStudy (anonymous):

No problem, I'm glad you speak it up so I know the right place to start :)

OpenStudy (anonymous):

The way you subtract equation is go vertically Pretending x is candy, y is cookie

OpenStudy (honeymonster):

so 3x -x? or x-3x?

OpenStudy (anonymous):

Either way is working, but in case you want 3x -x, then you need to switch the order of other terms, too!

OpenStudy (honeymonster):

i'll use x-3x then..would that mean -2x=-2

OpenStudy (honeymonster):

x=1 ???

OpenStudy (anonymous):

If you want (1) \ (2), then: x-3x = -2x, 5 -3 = 2 -> -2x = 2

OpenStudy (honeymonster):

yes of course..so x=-1???

OpenStudy (honeymonster):

ok...now how do we find y?

OpenStudy (anonymous):

Now, you has value x = -1, plug it back to any simpler equation => y

OpenStudy (honeymonster):

-1+2y=5 -1+5=6 6/2=3=y

OpenStudy (anonymous):

Yupe, so you have the intersection point of 2 lines I ( -1, 3) :D

OpenStudy (honeymonster):

thank you sooo much..your a life saver :)

OpenStudy (honeymonster):

makes me feel much more confident thanks

OpenStudy (honeymonster):

thank you :)

OpenStudy (anonymous):

Feel free to ask as many more as you like :)

OpenStudy (honeymonster):

i feel a bit cheeky but could you help me with this one please... 2x-4y=6 2x+y=8.5

OpenStudy (anonymous):

Just open the new post, I'll follow up :)

OpenStudy (anonymous):

Anyway since you notice both equations have the same 2x => Subtract them!

OpenStudy (honeymonster):

so 2x-2x? then 6-8.5 which would give x=-2.5?

OpenStudy (anonymous):

Just open the new post, I'll follow up :)